OpenStudy (anonymous):

express the following in simplest radical form. square root of 24

OpenStudy (anonymous):

@prettyboy20202

jimthompson5910 (jim_thompson5910):

gokart101 stop spamming, this isn't ebay

jimthompson5910 (jim_thompson5910):

The goal here is to factor 24 into two numbers where one is a perfect square, then we simplify from there \[\large \sqrt{24} = \sqrt{4*6}\] \[\large \sqrt{24} = \sqrt{4}*\sqrt{6}\] \[\large \sqrt{24} = 2\sqrt{6}\]